# Novel SiC UV Instrumentation Development with Potential Applications for the Habitable Worlds Observatory

### v1

@@
Prabal Saxena NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Corresponding author. Email: prabal.saxena@nasa.gov Zeynep Dilli NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Peter Snapp Detectors Systems Branch, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA Allison Youngblood Exoplanets and Stellar Astrophysics Lab, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA Tilak Hewagama NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Shahid Aslam NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Chullhee (Chace) Cho Detectors Systems Branch, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA Augustyn Waczynski Detectors Systems Branch, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA Nader Abuhassan SciGlob Instruments and Services, LLC., 9881 Broken Land Pkwy, Columbia, MD 21046, USA Ahn T. La NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Bryan K. Place SciGlob Instruments and Services, LLC., 9881 Broken Land Pkwy, Columbia, MD 21046, USA Thomas F. Hansico NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Ryan Stauffer NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Dina Bower NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Akin Akturk CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Neil Goldsman CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Bryce Galey CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Ethan Mountfort CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Mitchell Gross CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Ryan Purcell CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Usama Khalid CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Yekta Kamali CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Chris Darmody CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Robert Washington Howard University, 2400 6th St NW, Washington, DC 20059, USA Tim Livengood NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Daniel P. Moriarty NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Carl A. Kotecki Detectors Systems Branch, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA Narasimha S. Prasad NASA Langley Research Center, 1 Nasa Dr, Hampton, VA 23666, USA Joseph Wilkins Howard University, 2400 6th St NW, Washington, DC 20059, USA
@@ 1 Introduction
The most recent Astrophsysics Decadal Survey Report [1] underscored the importance of ultraviolet (UV) observations in exploring numerous high priority science questions. Ultraviolet observations are key to understanding stellar atmospheres, the morphology and dynamics of galaxies, the interaction of black holes with their surroundings and the properties of worlds outside our solar system, including their potential habitability. This is merely a subset of areas where UV observations are key, and such observations were also recognized as an essential means of achieving the objectives of the Report’s top flagship priority, the Habitable Worlds Observatory (HWO). The driving exoplanet and cosmological objectives of HWO require observations in the UV range of 100 – 320 nm using normal incidence optics [2]. These observations will require high-sensitivity UV detectors that can probe challenging observational targets at levels that require technological and engineering advancement. In this paper we detail recent and current work that is being carried out to fabricate and advance novel SiC UV instrumentation that is aimed at enabling more senstive measurements across numerous disciplines, with a short discussion of the promise such detectors may hold for HWO.
@@
The near-ultraviolet (NUV) spectral region (170-320 nm) is a subset of the ultraviolet spectral region that contains many powerful tracers of atoms, molecules, and small particulates such as dust and hazes. These NUV tracers have wide-ranging applications across numerous disciplines, such as Astrophysics, Planetary Science, Heliophysics, and Earth Science. Briefly, SiC detectors are a potentially powerful future technology in these other non-Astrophysics disciplines. With respect to Planetary Science, SiC detectors hold the potential to significantly improve measurements (1 order of magnitude improvement in signal-to-noise ratio (SNR) compared to state-of-the-art (SOA) silicon-based focal plane arrays) [3] for UV spectroscopy of comet coma and planetary atmosphere and aurora studies relevant to key planetary decadal objectives [4]. SiC detectors can also enable new measurements of molecules on bodies that are episodically sourced from plumes, for example, measurements of airglow emission when Io is in shadow (which could not historically be observed, due to the long accumulation time and contamination from scattered light from Jupiter) and could reveal the relative effects of SO2 frost and SO2 gas photoabsorption [5]. With respect to Heliophysics, SiC detectors have the potential to serve as radiation particle detectors in sun observations that require blindness to interfering visible and longer wavelength UV photons [6]. Lastly, applications specific to Earth Science include enabling ground/space-based passive remote sensing of trace gases and aerosols of interest by eliminating the need to conduct optical light filtering [7], which can interfere with accurate chemical retrievals by adding fringe structures to collected spectra that can obscure weaker chemical signatures. While SiC detectors have numerous other potential uses in all of these disciplines, this study focuses on technology development and application to an HWO astrophysics use case.
@@
With respect to Astrophysics, SiC detectors have the potential to provide a wide range of sensitive, enabling observations of interest, including UV spectroscopic measurements of active galactic nuclei, exoplanet host stars, and gravitational wave sources. SiC detectors directly satisfy the need for high-quantum efficiency (QE), solar-blind, broad-band NUV detectors identified in Astrophysics Biennial Technology Report 2022 [8]. In fact, the highest priority science areas from the Astrophsysics Decadal Survey Report [1] all benefit from sensitive NUV observations. For example, the NUV is a critical bandpass for distinguishing between physical models [9] of the extremely hot and rapidly evolving kilonova that emerges in the seconds to hours after the merger of two neutron stars. Additionally, in the search for habitable exoplanets, the Hartley-Huggins ozone band in the NUV is the most sensitive tracer of oxygenic atmospheres [10]. Obtaining information regarding these atmospheres is complicated by the high constrast ratios between the host star and planet and the low amount of reflected light from potential exo-Earth candidates. To enable significant breakthroughs in these areas, next generation detectors must maximize in-band sensitivity while minimizing noise. For imaging applications, minimizing responsivity to longer wavelength light is essential because many astrophysical sources are orders of magnitude brighter in the optical than the NUV. SiC is known to possess many of these properties [3] and with the advancement of fabrication techniques, is now being explored by our team as a potential option for many of these uses cases.
@@
Our team is pursuing fabrication of SiC instruments relevant to numerous disciplines under multiple projects funded by NASA’s Science Mission Directorate (SMD) and Space Technology Mission Directorate (STMD), as well as through Small Business Innovation Research Grants and through NASA Goddard’s Internal Research And Development program. The PReSSiC (Planetary Remote Sensing using SiC detectors) project is building, testing and validating a High Sensitivity Silicon Carbide Focal Plane Detector that would be optimized for miniaturized instrumentation that can be part of planetary science missions, including cubesats. The focal plane assembly is being designed so that it can obtain radiance observations with a higher detectivity than previous generations of NUV spectrometers from 200-340nm at a Rsimilar-$\sim$200. This is with the goal of being able to obtain remote sensing observations of targets in the NUV with a (SNR) $>$10$\times$ better than current leading detectors from identical observational platforms. Key science motivation in this project is understanding the hydration state of planetary bodies across the solar system, with a specific focus on the Moon and the identification of outgassed (non-water related) volatiles, characterization of atmospheric constituents and hazes and understanding of the mineralogy of airless bodies. Simultaneously, an ongoing development effort is the NASA STMD Early Career Initiative supported “Design of a Broad Utility, Visible-Blind, SiC, UV Spectrometer for Astrophysics, Planetary and Earth Science Applications, Validated Through Representative Measurement of Climate Forcing Gases” project, hereafter referred to as the Pandora-SiC effort, which aims to achieve clean differential optical absorption spectroscopy measurements of columnar trace gases in the earth atmosphere in the 280-450 nm regime through integration of visible blind UV detectors, produced by leveraging recent advances in SiC fabrication, into an existing, ground-based, atmospheric spectroscopy platform, Pandora [11]. SiC instrument development is also being carried out under other grants, including the Ultra-Violet Detector Innovation for Raman Exploration and CharacTerization (UV-DIRECT) of Ocean Worlds project. UV-DIRECT aims to enable identification of minerals, volatiles, organic molecules, biopolymers, water, and other hydrous phases in planetary materials by fabricating a SiC avalanche photodiode (APD), with high internal gain, specifically targeted for the identification of ocean world-relevant compounds. UV-DIRECT is being carried out in collaboration with Army Research Labs and is developing a detector that would use UV/NUV (266-340 nm) Raman spectroscopy with ppb sensitivity to achieve these planetary science observations. While a range of SiC instrumentation development is being carried out by our team, we will focus on the PReSSiC and Pandora-SiC plans and development in this paper and how advancements in these projects may be beneficial for future potential Astrophysics applications.
@@
Section 2 describes key background information on SiC as a semiconductor material, with a brief discussion of comparison to other materials. Section 3 lays out planned and current development work in the key SiC instrument development projects that are in progress. Section 4 provides updates on fabrication and integration progress of the photodectors for the corresponding projects. Section 5 discusses results and updates on recent and current testing of photodetector properties and also describes planned future work. Finally, Section 6 incorporates these SiC detector properties into a simulation of potential exoplanet observations by HWO using SiC detectors and also discusses potential application to HWO.
@@ 2 Background on SiC
For NASA, these wide energy band gap devices offered several key advantages for space applications over conventional Si ($E_{g}=1.1$ eV) based devices, such as:
  • visible-blind detection,
  • high thermal stability,
  • better radiation hardness,
  • high breakdown electric field,
  • high chemical inertness, and
  • greater mechanical strength.
Furthermore, the shorter cut-off wavelength of these material systems eliminates the need for bulky and expensive optical filtering components, mitigating risk and allowing for simpler optical designs of instrumentation.
A compelling reason for using wide-band gap semiconductors for UV detectors is the low dark currents that can be attained. The dark current, under zero or reverse bias conditions in the depletion region of a semiconductor, is determined from the rate of generation ($G$) of electron-hole pairs:
@@
where $q$ is the electronic charge, and $W$ is the depletion width. For silicon at room temperature, $n_{i}$ halves for every 9 K decrease in temperature. By comparison, a wider band gap semiconductor with the same density of states and $\tau$ as silicon, but a $0.036$ eV higher band gap, achieves the same factor of two reduction in $n_{i}$. The exponential dependence of $n_{i}$ on band gap makes a $0.12$ eV increase in band gap equivalent to a 30 K temperature drop, both offering an order-of-magnitude reduction in thermal generation rate. Wide-band gap semiconductors achieve significantly lower thermally generated dark current due to 10 or more orders of magnitude reduction in $n_{i}$, even in the presence of large differences in $\tau$. Furthermore, wide-band gap semiconductors exhibit strong temperature dependence, indicating that moderate cooling can significantly reduce the dark current.
Wide-band gap semiconductors are appealing for detecting UV-A (315–400 nm), UV-B (280–315 nm), UV-C (200–280 nm), and EUV (1–125 nm) radiation since they are inherently visible blind. The reduced need for blocking filters increases the system quantum efficiency (QE) and simplifies optical systems. Compared to narrow band gap materials, wide-band gap semiconductors are significantly more radiation tolerant, making them excellent candidates for space applications.
SiC detectors have several advantages over Si detectors:
@@
    Excellent Rigidity: SiC has high displacement energy ($\sim 21$ eV) compared to Si (12 eV), offering superior radiation hardness.
@@
Similar materials proposed for UV detection including indium gallium nitride (InGaN) [19], Diamond [20], and zinc oxide (ZnO) [21] suffer from similar lack of material maturity. With respect to another similar material, SiC may be able to provide improved measurements compared to current state of the art (SOA) delta-doped silicon CCDs utilized on missions like SPARCS ($>$4 times improved out of band rejection without a need for additional filters) [22]. Much of this is because SiC substrate and epi-growth technologies have developed to such a level as to allow the fabrication of many different types of SiC photodetectors with desired features [23]. SiC UV p-i-n photodiodes have already been fabricated and are commercially available. SiC avalanche photodiodes with extremely high gain and low excess noise have also been demonstrated. For additional information on the progress and history of SiC fabrication and use in many other applications, see [23].

### v2

@@
Prabal Saxena NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Corresponding author. Email: prabal.saxena@nasa.gov Zeynep Dilli CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Peter Snapp Detectors Systems Branch, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA Tilak Hewagama NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Shahid Aslam NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Chullhee Cho Cryogenics and Fluids Branch, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA Augustyn Waczynski Detectors Systems Branch, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA Nader Abuhassan SciGlob Instruments and Services, LLC., 9881 Broken Land Pkwy, Columbia, MD 21046, USA Anh T. La NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Bryan K. Place SciGlob Instruments and Services, LLC., 9881 Broken Land Pkwy, Columbia, MD 21046, USA Allison Youngblood NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Thomas F. Hansico NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Ryan Stauffer NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Dina Bower NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Akin Akturk CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Neil Goldsman CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Bryce Galey CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Ethan Mountfort CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Mitchell Gross CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Ryan Purcell CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Usama Khalid CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Yekta Kamali CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Chris Darmody CoolCAD Electronics, Inc., 5000 College Ave Suite 2105, College Park, MD 20740, USA Robert Washington Howard University, 2400 6th St NW, Washington, DC 20059, USA Tim Livengood NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Department of Astronomy, University of Maryland, College Park, MD 20740, USA Center for Research and Exploration in Space Science and Technology II Daniel P. Moriarty NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA Department of Astronomy, University of Maryland, College Park, MD 20740, USA Center for Research and Exploration in Space Science and Technology II Roser Juanola-Parramon NASA Goddard Space Flight Center, 8800 Greenbelt Rd, Greenbelt, MD 20771, USA University of Maryland, Baltimore County, 1000 Hilltop Cir, Baltimore, MD 21250 Carl A. Kotecki Detectors Systems Branch, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA Narasimha S. Prasad NASA Langley Research Center, 1 Nasa Dr, Hampton, VA 23666, USA Joseph Wilkins Howard University, 2400 6th St NW, Washington, DC 20059, USA
@@ 1 Introduction
The most recent Astrophysics Decadal Survey Report [1] underscored the importance of ultraviolet (UV) observations in exploring numerous high priority science questions. Ultraviolet observations are key to understanding stellar atmospheres, the morphology and dynamics of galaxies, the interaction of black holes with their surroundings and the properties of worlds outside our solar system, including their potential habitability. This is merely a subset of areas where UV observations are key, and such observations were also recognized as an essential means of achieving the objectives of the Decadal Report’s top flagship priority, the Habitable Worlds Observatory (HWO). The driving exoplanet and cosmological objectives of HWO require observations in the UV range of 100 – 320 nm using normal incidence optics [2]. These HWO will require UV detectors with sensitivity and noise properties significantly improved over the current state-of-the-art (SOA). In this paper we describe ongoing work to advance novel silicon carbide (SiC) detectors that would enable more sensitive UV measurements across numerous disciplines, with a short discussion of the promise such detectors may hold for HWO.
@@
The near-UV(NUV) spectral region (170-320 nm) is a subset of the UV spectral region that contains many powerful tracers of atoms, molecules, and small particulates such as dust and hazes. These NUV tracers have wide-ranging applications across numerous disciplines, such as Astrophysics, Planetary Science, Heliophysics, and Earth Science. Here we briefly describe science cases where SiC detectors would provide substantial gains over state-of-the-art silicon-based focal plane arrays [3]. A key advantage in all cases is SiC’s diminished sensitivity to visible light, which can often swamp desired UV signals. With respect to Planetary Science, order-of-magnitude signal-to-noise ratio improvements in UV spectroscopy of comet comae, planetary atmospheres, and aurorae would address key planetary decadal objectives [4]. SiC detectors can also enable new measurements of molecules on bodies that are episodically sourced from plumes, for example, measurements of airglow emission when Io is in shadow (which could not historically be observed, due to the long accumulation time and contamination from scattered light from Jupiter) and could reveal the relative effects of SO2 frost and SO2 gas photoabsorption [5]. Furthermore, improved SNR is critical for observing terrestrial bodies under challenging solar illumination conditions, such as observations of polar regions on the Moon. This is timely, as NASA’s Artemis and CLPS programs, as well as other international agencies, have targeted the lunar south polar region for human and robotic surface exploration within the next several years. Improved NUV observations of the lunar south pole has the ability to provide insight into latitude-dependent space weathering processes and probe compositional diversity and volatile distribution across the region at traverse-relevant spatial resolution, both providing context and guiding site selection and surface science activity prioritization [6, 7, 8]. With respect to Heliophysics, SiC detectors have the potential to serve as radiation particle detectors in solar observations that require blindness to interfering visible and longer wavelength UV photons[9]. Applications specific to Earth Science include ground- or space-based passive remote sensing of trace gases and aerosols. Optical light filtering is required for state-of-the-art Si detectors[10], and can interfere with accurate chemical retrievals by adding fringe structures to collected spectra that can obscure weaker chemical signatures. While SiC detectors have numerous other applications in each of these disciplines, this article focuses on technology development and application to an HWO astrophysics use case.
@@
With respect to Astrophysics, SiC detectors have the potential to provide a wide range of sensitive, enabling observations of interest, including UV spectroscopy of active galactic nuclei, exoplanet host stars, and gravitational wave sources. SiC detectors directly satisfy the need for high-quantum efficiency (QE), solar-blind, broad-band NUV detectors identified in Astrophysics Biennial Technology Report 2022 [11]. In fact, the highest priority science areas from the Astrophysics Decadal Survey Report [1] all benefit from sensitive NUV observations. For example, the NUV is a critical bandpass for distinguishing between physical models [12] of the extremely hot and rapidly evolving kilonova that emerges in the seconds to hours after the merger of two neutron stars. Additionally, in the search for habitable exoplanets, the Hartley-Huggins ozone band in the NUV is the most sensitive tracer of oxygenic atmospheres [13]. Obtaining information regarding these atmospheres is complicated by the high contrast ratios between the host star and planet and the low amount of reflected light from potential exo-Earth candidates. To enable significant breakthroughs in these areas, next generation detectors must maximize in-band sensitivity while minimizing noise. For imaging applications, minimizing responsivity to longer wavelength light is essential because many astrophysical sources are orders of magnitude brighter in the optical than the NUV. SiC possesses many of these properties [3] and with the advancement of fabrication techniques, is now being explored by our team as a potential option for many of these use cases.
@@
We are pursuing fabrication of SiC instruments relevant to numerous disciplines under multiple projects funded by NASA’s Science Mission Directorate (SMD) and Space Technology Mission Directorate (STMD), as well as through Small Business Innovation Research Grants and NASA Goddard Space Flight Center’s Internal Research and Development program. The PReSSiC (Planetary Remote Sensing using SiC detectors - PIs Saxena/Hewagama) project [14] is building, testing and validating a High Sensitivity Silicon Carbide Focal Plane Detector optimized for miniaturized instrumentation that can be part of remote sensing planetary science missions, including cubesats. The focal plane assembly would obtain radiance observations with a higher detectivity than previous generations of NUV spectrometers from 200-340nm at resolving power Rsimilar-$\sim$200, enabling SNR $>$10$\times$ better than current leading detectors from identical observational platforms. The key science motivation for PReSSiC is understanding the hydration state of planetary bodies across the solar system, with a specific focus on the Moon and the identification of outgassed (non-water related) volatiles, characterization of atmospheric constituents and hazes and understanding of the mineralogy of airless bodies. Simultaneously, an ongoing development effort is the NASA STMD Early Career Initiative supported “Design of a Broad Utility, Visible-Blind, SiC, UV Spectrometer for Astrophysics, Planetary and Earth Science Applications, Validated Through Representative Measurement of Climate Forcing Gases” project (PI Snapp), hereafter referred to as the Pandora+SiC project. Pandora+SiC aims to achieve clean differential optical absorption spectroscopy of columnar trace gases in Earth’s atmosphere in the 280-450 nm range through integration of visible-blind UV detectors, produced by leveraging recent advances in SiC fabrication, into an existing, ground-based, atmospheric spectroscopy platform, Pandora [15]. SiC instrument development is also being carried out under another NASA ROSES PICASSO award, namely, the Ultra-Violet Detector Innovation for Raman Exploration and CharacTerization (UV-DIRECT) of Ocean Worlds project (PI Bower). UV-DIRECT aims to enable identification of minerals, volatiles, organic molecules, biopolymers, water, and other hydrous phases in planetary materials by fabricating a SiC avalanche photodiode (APD), with high internal gain, specifically targeted for applications to ocean worlds surface missions that require robust and compound-specific instrumentation. UV-DIRECT aims to enable identification of minerals, volatiles, organic molecules, biopolymers, water, and other hydrous phases in planetary materials by fabricating a SiC avalanche photodiode (APD), with high internal gain, specifically targeted for the identification of ocean world-relevant compounds [16]. UV-DIRECT is being carried out in collaboration with Army Research Labs and is developing a detector that would use UV/NUV (266-340 nm) Raman spectroscopy with ppb sensitivity to achieve these planetary science observations [17]. Raman spectroscopy is mainly an in-situ technique, but since SiC APD technology is so versatile, the development efforts and results for UV-DIRECT have fed into the other projects more relevant to HWO [18]. While a range of SiC instrumentation development is being carried out by our team, we will focus on the PReSSiC and Pandora+SiC plans and development in this paper and how advancements in these projects may be beneficial for future potential Astrophysics applications.
@@
Section 2 describes key background information on SiC as a semiconductor material, with a brief discussion of comparison to other materials. Section 3 lays out planned and current development work in the key SiC instrument development projects that are in progress. Section 4 provides updates on fabrication and integration progress of the photodectors for the corresponding projects. Section 5 discusses results and updates on recent and current testing of photodetector properties as well as planned future work. Finally, Section 6 incorporates these SiC detector properties into a simulation of exo-Earth observations by HWO.
@@ 2 Background on SiC
For NASA, these wide energy band gap devices offered several key advantages for space applications over conventional Si ($E_{g}=1.1$ eV) based devices. These include visible-blind detection, high thermal stability, better radiation hardness, high breakdown electric field, high chemical inertness, and greater mechanical strength. Additional detail on these properties is given later in this section. Furthermore, the shorter cut-off wavelength of these material systems eliminates the need for bulky and expensive optical filtering components, mitigating risk and allowing for simpler optical designs, and most importantly increasing in-band throughput. .
The most compelling reason for using wide-band gap semiconductors for UV detectors is the low dark currents that can be attained. The dark current, under zero or reverse bias conditions in the depletion region of a semiconductor, is determined from the rate of generation ($G$) of electron-hole pairs:
@@
where $q$ is the electronic charge, and $W$ is the depletion width. For silicon at room temperature, $n_{i}$ halves for every 5.5 K decrease in temperature [25]. By comparison, a wider band gap semiconductor with the same density of states and $\tau$ as silicon, but a $0.036$ eV higher band gap, achieves the same factor of two reduction in $n_{i}$. The exponential dependence of $n_{i}$ on band gap makes a $0.12$ eV increase in band gap equivalent to a 30 K temperature drop, both offering an order-of-magnitude reduction in thermal generation rate. Wide-band gap semiconductors achieve significantly lower thermally generated dark current due to 10 or more orders of magnitude reduction in $n_{i}$, even in the presence of large differences in $\tau$, and therefore have applications in high temperature space environments. Furthermore, wide-band gap semiconductors exhibit strong temperature dependence, indicating that moderate cooling can significantly reduce the dark current.
Wide-band gap semiconductors have additional benefits for detecting UV-A (315–400 nm), UV-B (280–315 nm), UV-C (200–280 nm), and EUV (1–125 nm) radiation since they are inherently visible blind. This visible blindness significantly reduces the need for blocking filters which simplifies optical system design and reduces the size, weight, power consumption , and cost (SWaP-C) of UV instrument, and most importantly increases the system quantum efficiency (QE) allowing for more senitive measurements
Compared to narrow band gap materials, wide-band gap semiconductors are significantly more radiation tolerant, making them excellent candidates for space applications. SiC offers superior radiation hardness (long-term resilience) compared to conventional Si due to its stable crystal structure resulting from stronger covalent bonds than Si, and thereby higher displacement threshold with low defect mobility, which reduces radiation-induced damage. SiC’s stronger covalent bonds and higher melting point make it more resilient to radiation exposure, as it can maintain its structural integrity even under extreme condition like high-radiation environment of space. When SiC gets cooled to lower temperature, SiC’s radiation resistance can be further enhanced owing to the reduced thermal vibrations and lower defect mobility preventing defects migrating or clustering. Moreover, the reduction in charge carrier mobility at lower temperature prevents radiation-induced disruptions in SiC’s electrical performance, making it an ideal candidate for space applications.
In summary, SiC detectors have several advantages over Si detectors:
@@
    Excellent Radiation Hardness: SiC has high displacement energy ($\sim 21$ eV) compared to Si (12 eV), offering superior radiation hardness.